# A Simple LLM Framework for Long-Range Video Question-Answering

### v1

@@ Abstract
We present LLoVi, a language-based framework for long-range video question-answering (LVQA). Unlike prior long-range video understanding methods, which are often costly and require specialized long-range video modeling design (e.g., memory queues, state-space layers, etc.), our approach uses a frame/clip-level visual captioner (e.g., BLIP2, LaViLa, LLaVA) coupled with a Large Language Model (GPT-3.5, GPT-4) leading to a simple yet surprisingly effective LVQA framework. Specifically, we decompose short and long-range modeling aspects of LVQA into two stages. First, we use a short-term visual captioner to generate textual descriptions of short video clips (0.5-8s in length) densely sampled from a long input video. Afterward, an LLM aggregates the densely extracted short-term captions to perform long-range temporal reasoning needed to understand the whole video and answer a question. To analyze what makes our simple framework so effective, we thoroughly evaluate various components of our system. Our empirical analysis reveals that the choice of the visual captioner and LLM is critical for good LVQA performance. Furthermore, we show that a specialized prompt that asks the LLM first to summarize the noisy short-term visual captions and then answer a given input question leads to a significant LVQA performance boost. On EgoSchema, which is best known as a very long-form video question-answering benchmark, our method achieves 50.3% accuracy, outperforming the previous best-performing approach by 18.1% (absolute gain). In addition, our approach outperforms the previous state-of-the-art by 4.1% and 3.1% on NeXT-QA and IntentQA. We also extend LLoVi to grounded LVQA and show that it outperforms all prior methods on the NeXT-GQA dataset. We will release our code at https://github.com/CeeZh/LLoVi.
@@ 1 Introduction
Figure 1: Comparison of long-range video question-answering (LVQA) accuracy on EgoSchema, the new LVQA benchmark that addresses many limitations of prior long-range video benchmarks. Our simple LLM framework, named LLoVi, outperforms all prior LVQA approaches by a significant margin (+18.1%).
Recent years have witnessed remarkable progress in short video understanding (5-15s in length) [61, 77, 11, 73, 63]. However, extending these models to long videos (e.g., several minutes or hours in length) is not trivial due to the need for sophisticated long-range temporal reasoning capabilities. Most existing long-range video models rely on costly and complex long-range temporal modeling schemes, which include memory queues [67, 4, 24, 23], long-range feature banks [66, 5, 82], space-time graphs [18, 60], state-space layers [19, 20, 56] and other complex long-range modeling modules [17, 1, 76].
Recently, Large Language Models (LLMs) have shown impressive capability for long-range reasoning on a wide range of tasks such as document understanding [49, 58, 14] and long-horizon planning [35, 15, 47]. Motivated by these results in the natural language and decision-making domain, we explore using LLMs for long-range video question answering (LVQA). Specifically, we propose LLoVi, a simple language-based framework for long-range video understanding. Unlike prior long-range video models, our approach does not require specialized long-range video modules (e.g., memory queues, state-space layers, spatiotemporal graphs, etc.) but instead uses a short-term visual captioner coupled with an LLM, thus exploiting the long-range temporal reasoning ability of LLMs. Our simple two-stage approach tackles the LVQA task by decomposing it into short and long-range modeling subproblems:
@@
    First, given a long video input, we segment it into multiple short clips and convert them into one-sentence textual descriptions using a pre-trained frame/clip-level visual captioner (e.g., BLIP2, LaViLa).
@@
    Afterward, we concatenate the temporally ordered captions from Step 1 and feed them into an LLM (e.g., GPT-3.5, GPT-4) with proper prompts to perform long-range video reasoning for LVQA.
Recent works have explored incorporating visual captioning models into LLMs for video understanding. Several methods proposed using LLMs for various short-term video understanding tasks. For example, Video ChatCaptioner [3] generates enriched video descriptions by learning to answer questions from ChatGPT. Similarly, ChatVideo [55] stores the detected tracklets in a database, which are then used for interacting with the user. Additionally, Socratic Models [81], VidIL [62], and VideoChat [31] all use pretrained visual models to extract low-level video concepts (e.g., objects, actions, scenes, etc.) and then leverage LLM to perform various short-term video understanding tasks. Beyond short-term video understanding tasks, we note that the concurrent work in [10] applies an LLM-based framework for long-range video understanding. However, their analysis is limited to movie-based datasets that rely heavily on non-visual inputs such as speech and subtitles, thus requiring limited visual analysis [37]. Compared to this concurrent work, we conduct our experiments on EgoSchema, the newly introduced LVQA benchmark that addresses many limitations of prior long-range video benchmarks. Furthermore, unlike prior LLM-based video understanding frameworks, many of which perform only qualitative analysis of their models [3, 55, 31, 33], we conduct a thorough empirical study investigating the effectiveness of our framework.
Figure 2: An illustration of LLoVi, our simple LLM framework for long-range video question-answering (LVQA). We use Large Language Models (LLMs) like GPT-3.5 and GPT-4 for their long-range modeling capabilities. Our method involves two stages: first, we use short-term visual captioners (e.g., LaViLa, BLIP2) to generate textual descriptions for brief video clips (0.5s-8s). Then, an LLM aggregates these dense, short-term captions for long-range reasoning required for LVQA. This simple approach yields impressive results, demonstrating LLMs’ effectiveness in LVQA.
@@
Specifically, we investigate (i) the selection of a visual captioner, (ii) the choice of an LLM, (iii) the LLM prompt design, (iv) few-shot in-context learning, (v) optimal video processing configurations (i.e., clip length, sampling rate, etc.), and (vi) the generalization of our framework to other datasets and tasks. Our key empirical findings include:
  • A multi-round prompt that first asks the LLM to summarize the noisy/redundant short-term visual captions and then answer a given question based on the summarized captions leads to the most significant boost in performance (+5.8%) among the prompts we have tried (e.g., zero-shot CoT, Self-Consistency).
  • GPT-3.5 provides the best tradeoff between accuracy, computational cost, and sufficiently large context length.
@@
  • Using LaViLa visual captioner [83] leads to best results (51.8%) followed by BLIP-2 [28] (46.7%) and EgoVLP [43] (46.6%).
  • Few-shot in-context learning leads to a large improvement on both the variant of our model with the simplest prompt (+4.7%) and our best-performing variant with a multi-round prompt (+4.1%).
@@
  • Densely extracting visual captions from consecutive 1-second video clips of the long video input leads to the strongest performance on EgoSchema.
  • Our final method, built using the above-listed empirical insights, achieves 50.3% zero-shot LVQA accuracy on the full test set of EgoSchema, outperforming the previous best approach by 18.1% as shown in Figure 1.
  • Our LLoVi also outperforms prior approaches on NeXT-QA and IntentQA by 4.1% and 3.1%, and it also achieves state-of-the-art performance on NeXT-GQA, a recently introduced grounded LVQA benchmark.
We hope that our simple, training-free method will encourage new ideas and a simpler model design in LVQA. We will release our code to enable the community to build on our work.
@@ 2 Related Work
Long-range Video Understanding. Modeling long-range videos (e.g., several minutes or longer) typically requires models with sophisticated temporal modeling capabilities, often leading to complex model design. LF-VILA [50] proposes a Temporal Window Attention (HTWA) mechanism to capture long-range dependency in long-form video. MeMViT [67] and MovieChat [48] adopt a memory-based design to store information from previously processed video segments. Several prior methods use space-time graphs [18, 60] or relational space-time modules [76] to capture spatiotemporal dependencies in long videos. Lastly, the recently introduced S4ND [40], ViS4mer [19] and S5 [56] use Structured State-Space Sequence (S4) [13] layers to capture long-range dependencies in the video. Unlike these prior approaches, we do not propose any complex long-range temporal modeling modules but instead develop a simple and strong LLM-based framework for zero-shot LVQA.
@@
LLMs for Video Understanding. The recent surge in large language models (LLMs) [2, 41, 54, 46, 9, 53] has inspired many LLM-based applications in video understanding. Models like Socratic Models [81] and VideoChat [31] integrate pretrained visual models with LLMs for extracting visual concepts and applying them to video tasks. Video ChatCaptioner [3] and ChatVideo [55] leverage LLMs for video representation and dialog-based user interaction, respectively. VidIL [62] employs LLMs for adapting image-level models to video tasks using few-shot learning. Beyond short-term understanding, studies [33, 10] have explored LLMs for long-range video modeling. The work in [33] uses GPT-4 for video summarization but lacks quantitative evaluation. Meanwhile, [10] focuses on movie datasets, requiring limited visual analysis [37]. In contrast, we conduct our experiments on the EgoSchema benchmark for long-range LVQA and provide an extensive empirical analysis of various design choices of our simple approach.
@@
Video Question Answering. Unlike image question-answering, video question-answering (VidQA) presents unique challenges, requiring both spatial and temporal reasoning. Most existing VidQA methods, either using pretraining-finetuning paradigms [6, 26, 79], zero-shot [74, 51, 34, 79], or few-shot learning [62], focus on short-term video analysis (5-30s). To overcome the limitations of short-term VidQA, new benchmarks have been proposed covering longer video durations: NextQA [68] averages 44s, while ActivityNet-QA [80], TVQA [25], How2QA [72], MovieQA [52], and DramaQA [7] range from 100s to several minutes. Despite these lengths, works in  [37, 75, 21] found that many benchmarks can be solved by analyzing only brief clips, not requiring extensive video modeling, and exhibited language biases, being solvable using pure text-only methods that ignore visual content. To address these issues, the EgoSchema benchmark [37] was recently introduced, requiring at least 100 seconds of video analysis and not exhibiting any language biases. Thus, our work focuses on the LVQA analysis on EgoSchema.
@@ 3 Method
Figure 3: An illustration of the multi-round summarization-based prompt. Our empirical analysis shows that many modern LLMs may struggle when provided with long ($>$1K words), noisy, and redundant/irrelevant caption sequences. To address this issue, we study a more specialized LLM prompt that asks an LLM first to summarize the noisy short-term visual captions (first round of prompting) and then answer a given question about the video (second round of prompting). Our results indicate that such a multi-round prompting strategy significantly boosts LVQA performance (+5.8%) compared to standard prompting techniques.
Recently, LLMs have been shown to excel on a wide range of long-range modeling tasks [35, 15, 47, 49, 58, 14]. Motivated by these long-range modeling capabilities, we propose to tackle the long-range video question-answering (LVQA) task by decomposing it into two subtasks: 1) short-term video clip captioning, and 2) long-range text-based video understanding. Our motivation for such decomposition is to separate the short-term and long-range video modeling so that we could leverage the strong existing short-term visual captioners (e.g., LaViLa, BLIP2) for the first sub-task, and powerful zero-shot LLMs (e.g., GPT-3.5, GPT-4, LLaMA, T5) for the second sub-task, thus, exploiting the power of LLMs for long-range modeling.
Our decomposed LVQA framework, named LLoVi (Language-based Long-range Video Question-answering), brings important advantages. First, our approach is simple as it does not rely on complex/specialized long-range video modeling operators (e.g., memory queues, state-space layers, space-time graphs, etc.). Second, our framework is training-free, which makes it easy to apply it to LVQA in zero-shot settings. Third, our method is agnostic to the exact choices of a visual captioner and LLM, and thus, it can benefit from future improvements in visual captioning/LLM model design. Figure 2 presents a detailed illustration of our high-level approach. Below, we provide details about each component of our framework.
@@ 3.1 Short-term Video Clip Captioning
Given a long untrimmed video input $V$, we first segment it into $N_{v}$ non-overlapping short video clips $v=\{v_{m}\}_{m=1}^{N_{v}}$, where $v_{m}\in\mathbb{R}^{T_{v}\times H\times W\times 3}$ and $T_{v},H,W$ are the number of frames, height and width of a short video clip respectively. Afterward, we feed each video clip $v_{m}$ into a pretrained short-term visual captioner $\phi$, which produces textual captions $c_{m}=\phi(v_{m})$, where $c_{m}=(w_{1},\ldots,w_{L_{m}})$ and $w_{i}$ represents the i-th word in caption $c_{m}$ of length $L_{m}$. Note that our model is not restricted to any specific visual captioning model. Our experimental section demonstrates that we can incorporate various video (LaViLa [83], EgoVLP [43], VideoBLIP [78]) and image (BLIP-2 [27]) captioning models. Next, we describe how our extracted short-term captions are processed by an LLM.
@@ 3.2 Long-range Reasoning with an LLM
We want to leverage foundational LLMs for holistic long-range video understanding rather than relying on complex long-range temporal modeling schemes (e.g., memory queues, state-space layers, space-time graphs, etc.) used by prior methods  [61, 67].
Formally, given short-term visual captions $\{c_{m}\}_{m=1}^{N_{v}}$ for all $N_{v}$ short video clips, we first concatenate the clip captions into the full video captions $C=[c_{1},\ldots,c_{N_{v}}]$ in the same order as the captions appear in the original video. Afterward, the concatenated video captions $C$ are fed into an LLM for long-range video reasoning. Specifically, given the concatenated video captions $C$, the question $Q$, and the answer candidates $A$, we prompt the LLM to select the correct answer using the following prompt template: “Please provide a single-letter answer (A, B, C, D, E) to the following multiple-choice question $\{Q\}$. You are given language descriptions of a video. Here are the descriptions: $\{C\}$. Here are the choices $\{A\}$.”. The full prompt is included in Supplementary Materials.
Our experiments in Section 4.3 suggest that this simple approach works surprisingly well for LVQA. However, we also discovered that many modern LLMs (e.g., GPT-3.5, LLaMA) may struggle when provided with long ($>$1K words), noisy, and potentially redundant/irrelevant caption sequences. To address these issues, we investigate more specialized LLM prompts that ask an LLM first to summarize the noisy short-term visual captions (first round of prompting) and then answer a given question about the video (second round of prompting). Specifically, we formulate such a multi-round prompt as follows: given the video captions $C$, the question $Q$, and the answer candidates $A$, instead of directly feeding the $\{C,Q,A\}$ triplet into LLM for LVQA, we first ask the LLM to provide a summary of the captions in the first round, which we denote as $S$ using the following prompt template: “You are given language descriptions of a video: {$C$}. Please give me a {$N_{w}$} word summary.$N_{w}$ denotes the desired number of words in the summary $S$. Afterward, during the second round of prompting, instead of using the captions $C$, we use the summary $S$ as input for the LLM to select one of the answer candidates. Conceptually, this may be beneficial, as the LLM-generated summary $S$ filters out potentially irrelevant/noisy information from the initial set of captions $C$, making LLM inputs for the subsequent QA process more succinct and cleaner. A detailed illustration of our multi-round prompt is shown in Figure 3.

### v3

@@ Abstract
We present LLoVi, a simple yet effective Language-based Long-range Video question-answering (LVQA) framework. Our method decomposes the short- and long-range modeling aspects of LVQA into two stages. First, we use a short-term visual captioner to generate textual descriptions of short video clips (0.5-8 seconds in length) densely sampled from a long input video. Afterward, an LLM aggregates the densely extracted short-term captions to answer a given question. Furthermore, we propose a novel multi-round summarization prompt that asks the LLM first to summarize the noisy short-term visual captions and then answer a given input question. To analyze what makes our simple framework so effective, we thoroughly evaluate various components of our framework. Our empirical analysis reveals that the choice of the visual captioner and LLM is critical for good LVQA performance. The proposed multi-round summarization prompt also leads to a significant LVQA performance boost. Our method achieves the best-reported results on the EgoSchema dataset, best known for very long-form video question-answering. LLoVi also outperforms the previous state-of-the-art by 10.2% and 6.2% on NExT-QA and IntentQA for LVQA. Finally, we extend LLoVi to grounded VideoQA, which requires both QA and temporal localization, and show that it outperforms all prior methods on NExT-GQA. Code is available at https://github.com/CeeZh/LLoVi.
@@ 1 Introduction
Figure 1: Comparison between LLoVi (ours) and the recent FrozenBiLM Yang et al. (2022a) video QA method. Like most prior methods, FrozenBiLM is best suited for short-range video understanding. Thus, as illustrated in the figure, it fails to answer a question that requires reasoning about complex human activities in a long video. In comparison, our method effectively reasons over long temporal extents and produces a correct answer.
Recent years have witnessed remarkable progress in short video understanding (5-15s in length) Wang et al. (2022a); Ye et al. (2023); Fu et al. (2021); Yang et al. (2022a); Wang et al. (2023g). However, extending these models to long videos (e.g., several minutes or hours in length) is not trivial due to the need for sophisticated long-range temporal reasoning capabilities. Most existing long-range video models rely on costly and complex long-range temporal modeling schemes, which include memory queues Wu et al. (2022); Chen et al. (2020); Lee et al. (2021, 2018), long-range feature banks Wu et al. (2019); Cheng and Bertasius (2022); Zhang et al. (2021), space-time graphs Hussein et al. (2019b); Wang et al. (2021), state-space layers Islam and Bertasius (2022); Islam et al. (2023); Wang et al. (2023a) and other complex long-range modeling modules Hussein et al. (2019a); Bertasius et al. (2021); Yang et al. (2023).
Recently, Large Language Models (LLMs) have shown impressive capability for long-range reasoning on a wide range of tasks such as document understanding Sun et al. (2023); Wang et al. (2023e); Gur et al. (2023) and long-horizon planning Liu et al. (2023a); Hao et al. (2023); Song et al. (2023a). Motivated by these results in the natural language and decision-making domain, we explore using LLMs for long-range video question answering (LVQA). Specifically, we propose LLoVi, a simple yet effective language-based framework for long-range video understanding. Unlike prior long-range video models, our approach does not require specialized long-range video modules (e.g., memory queues, state-space layers, etc.) but instead uses a short-term visual captioner coupled with an LLM, thus exploiting the long-range temporal reasoning ability of LLMs. Our simple two-stage framework tackles the LVQA task by decomposing it into short and long-range modeling subproblems:
@@
    First, given a long video input, we segment it into multiple short clips and convert them into short textual descriptions using a pre-trained frame/clip-level visual captioner (e.g., BLIP2 Li et al. (2023c), LaViLa Zhao et al. (2023), LLaVA Liu et al. (2023b)).
@@
    Afterwards, we concatenate the temporally ordered captions from Step 1 and feed them into an LLM (e.g., GPT-3.5, GPT-4, LLaMA) to perform long-range reasoning for LVQA.
To further enhance the effectiveness of our framework, we also introduce a novel multi-round summarization prompt that asks the LLM first to summarize the short-term visual captions and then answer a given question based on the LLM-generated video summary. Since the generated captions may be noisy or redundant, such a summarization scheme enables filtering out potentially distracting/irrelevant information and eliminating redundant sentences, which significantly improves the reasoning ability of the LLM for LVQA.
@@
Additionally, we conduct an empirical study on EgoSchema to investigate the factors behind our framework’s success. Specifically, we study (i) the selection of a visual captioner, (ii) the choice of an LLM, (iii) the LLM prompt design, and (iv) optimal video processing configurations. Our key empirical findings include:
  • Our newly proposed multi-round summarization prompt leads to the most significant boost in performance (+3.6%) among the prompts we have tried (e.g., Chain-of-Thought, Plan-and-Solve).
  • GPT-4 as an LLM provides the best performance, while GPT-3.5 provides the best trade-off between the accuracy and the cost.
@@
  • LaViLa Zhao et al. (2023) as a visual captioner produces best results (55.2%) followed by BLIP-2 Li et al. (2023c) (50.6%) and EgoVLP Qinghong Lin et al. (2022) (46.6%).
@@
  • Extracting visual captions from consecutive 1-second video clips of the long video input leads to the best results. Also, extracting captions from sparsely sampled video clips leads to 8x improved efficiency while still maintaining reasonable performance (2.0% accuracy drop).
We want to make it clear that LLoVi is not based on any complex or novel design choices. It is a simple, effective, and training-free method that outperforms all prior approaches on EgoSchema, NExT-QA, IntentQA, and NeXT-GQA, establishing a strong baseline for the LVQA task. We hope that our work will encourage the LVQA community to build on our work and use our thorough empirical insights to develop new LVQA models.
@@ 2 Related Work
Long-range Video Understanding. Modeling long-range videos (e.g., several minutes or longer) typically requires models with sophisticated temporal modeling capabilities, often leading to complex model design. LF-VILA Sun et al. (2022) proposes a Temporal Window Attention (HTWA) mechanism to capture long-range dependency in long-form video. MeMViT Wu et al. (2022) and MovieChat Song et al. (2023b) adopt a memory-based design to store information from previously processed video segments. Several prior methods use space-time graphs Hussein et al. (2019b); Wang et al. (2021) or relational space-time modules Yang et al. (2023) to capture spatiotemporal dependencies in long videos. Lastly, the recently introduced S4ND Nguyen et al. (2022), ViS4mer Islam and Bertasius (2022) and S5 Wang et al. (2023a) use Structured State-Space Sequence (S4) Gu et al. (2021) layers to capture long-range dependencies in the video. Unlike these prior approaches, we do not use any complex long-range temporal modeling modules but instead develop a simple and strong LLM-based framework for zero-shot LVQA.
@@
LLMs for Video Understanding. The recent surge in large language models (LLMs) Brown et al. (2020); OpenAI (2023b); Touvron et al. (2023); Raffel et al. (2020); Chung et al. (2022); Tay et al. (2022) has inspired many LLM-based applications in video understanding. Methods like Socratic Models Zeng et al. (2022) and VideoChat Li et al. (2023e) integrate pretrained visual models with LLMs for extracting visual concepts and applying them to video tasks. Video ChatCaptioner Chen et al. (2023) and ChatVideo Wang et al. (2023b) leverage LLMs for video representation and dialog-based user interaction, respectively. VidIL Wang et al. (2022b) employs LLMs for adapting image-level models to video tasks using few-shot learning. Beyond short-term video understanding, the works in Lin et al. (2023a); Chung and Yu (2023); Bhattacharya et al. (2023) explored LLMs for long-range video modeling. The work in Lin et al. (2023a) uses GPT-4 for various long-range video modeling tasks but lacks quantitative evaluation. Meanwhile, Chung and Yu (2023) focuses on movie datasets, requiring limited visual analysis Mangalam et al. (2023) and mostly relying on non-visual speech/subtitle inputs. In contrast to these prior methods, we focus on the LVQA task and provide an extensive empirical analysis of various design choices behind our LLM framework.
@@
Video Question Answering. Unlike image question-answering, video question-answering (VidQA) presents unique challenges, requiring both spatial and temporal reasoning. Most existing VidQA methods, either using pretraining-finetuning paradigms Cheng et al. (2023); Lei et al. (2021); Yu et al. (2023), zero-shot Yang et al. (2022b); Surís et al. (2023); Lin et al. (2023b); Yu et al. (2023), or few-shot learning Wang et al. (2022b), focus on short-term video analysis (5-30s). To overcome the limitations of short-term VidQA, new benchmarks have been proposed: ActivityNet-QA Yu et al. (2019), TVQA Lei et al. (2018), How2QA Yang et al. (2021), MovieQA Tapaswi et al. (2016), and DramaQA Choi et al. (2021) ranging from 100s to several minutes in video duration. Despite longer video lengths, the analysis in  Mangalam et al. (2023); Yang et al. (2020); Jasani et al. (2019) found that many of these benchmarks can be solved by analyzing only short clips (i.e., not requiring long-range video modeling) or by using pure text-only methods that ignore visual content. To address these issues, the EgoSchema benchmark Mangalam et al. (2023) was recently introduced, requiring at least 100 seconds of video analysis and not exhibiting language-based biases.
Figure 2: An illustration of LLoVi, our simple LLM framework for long-range video question-answering (LVQA). We use Large Language Models (LLMs) like GPT-3.5 and GPT-4 for their long-range modeling capabilities. Our method involves two stages: first, we use short-term visual captioners (e.g, LaViLa, BLIP2) to generate textual descriptions for brief video clips (0.5s-8s). Then, an LLM aggregates these dense, short-term captions for long-range reasoning required for LVQA. This simple approach yields impressive results, demonstrating LLMs’ effectiveness in LVQA.
LLM Prompt Design. With the emergence of LLMs, there has been an increasing research emphasis on LLM prompt design. The recent works in Wei et al. (2022); Zhou et al. (2023); Schick and Schütze (2020); Chen et al. (2022); Yao et al. (2022) explored prompting strategy in few-shot learning settings. To eliminate the need for extensive human annotations, Kojima et al. (2022); Wang et al. (2023c, f) proposed zero-shot prompting methods. Subsequent research Zhou et al. (2022); Zhang et al. (2022); Pryzant et al. (2023) has concentrated on the automatic refinement of prompts. Instead, we propose a multi-round summarization LLM prompt for handling long, noisy, and redundant textual inputs describing video content for LVQA.
@@ 3 Method
Figure 3: An illustration of our multi-round summarization prompt that first asks an LLM to summarize the noisy short-term visual captions (first round of prompting) and then answer a given question about the video based on the LLM-generated summary (second round of prompting). Our results indicate that such a multi-round prompting strategy significantly boosts LVQA performance compared to standard prompting techniques (+5.8%).
Our method, LLoVi, decomposes LVQA into two subtasks: 1) short-term video clip captioning and 2) long-range text-based video understanding. Our decomposed LVQA framework brings several important advantages. First, our approach is simple as it does not rely on complex/specialized long-range video modeling operators (e.g., memory queues, state-space layers, space-time graphs, etc.). Second, our framework is training-free, which makes it easy to apply it to LVQA in zero-shot settings. Third, our framework enables us to leverage the strong existing short-term visual captioners (e.g., LaViLa, LLaVA) and powerful zero-shot LLMs (e.g., GPT-3.5, GPT-4, LLaMA). Fourth, our method is highly flexible, i.e., it can incorporate various visual captioners and LLMs, and also benefit from future improvements in visual captioning/LLM model design. Figure 2 presents a detailed illustration of our high-level approach. Below, we provide details about each component of our framework.
@@ 3.1 Short-term Video Clip Captioning
Given a long untrimmed video input $V$, we first segment it into $N_{v}$ non-overlapping short video clips $v=\{v_{m}\}_{m=1}^{N_{v}}$, where $v_{m}\in\mathbb{R}^{T_{v}\times H\times W\times 3}$ and $T_{v},H,W$ are the number of frames, height and width of a short video clip respectively. Afterward, we feed each video clip $v_{m}$ into a pretrained short-term visual captioner $\phi$, which produces textual captions $c_{m}=\phi(v_{m})$, where $c_{m}=(w_{1},\ldots,w_{L_{m}})$ and $w_{i}$ represents the i-th word in caption $c_{m}$ of length $L_{m}$. Note that our model is not restricted to any specific visual captioning model. Our experimental section demonstrates that we can incorporate various video (LaViLa Zhao et al. (2023), EgoVLP Qinghong Lin et al. (2022), and image (BLIP-2 Li et al. (2023d)) captioning models. Next, we describe how our extracted short-term captions are processed by an LLM.
@@ 3.2 Long-range Reasoning with an LLM
We want to leverage foundational LLMs for holistic long-range video understanding. Formally, given short-term visual captions $\{c_{m}\}_{m=1}^{N_{v}}$ for all $N_{v}$ short video clips, we first concatenate the clip captions into the full video captions $C=[c_{1},\ldots,c_{N_{v}}]$ in the same order as the captions appear in the original video. Afterward, the concatenated video captions $C$ are fed into an LLM for long-range video reasoning. Specifically, given the concatenated video captions $C$, the question $Q$, and the answer candidates $A$, we prompt the LLM to select the correct answer using the following prompt template: “Please provide a single-letter answer (A, B, C, D, E) to the following multiple-choice question $\{Q\}$. You are given language descriptions of a video. Here are the descriptions: $\{C\}$. Here are the choices $\{A\}$.". The full prompt is included in the Supplementary Material.
Our experiments in Section 4.3 suggest that this simple approach works surprisingly well for LVQA. However, we also discovered that many modern LLMs (e.g., GPT-3.5, LLaMA) may struggle when provided with long ($>$1K words), noisy, and potentially redundant/irrelevant caption sequences. To address these issues, we investigate more specialized LLM prompts that ask an LLM first to summarize the noisy short-term visual captions (first round of prompting) and then answer a given question about the video (second round of prompting). Specifically, we formulate such a multi-round prompt as follows: given the video captions $C$, the question $Q$, and the answer candidates $A$, instead of directly feeding the $\{C,Q,A\}$ triplet into LLM for LVQA, we first ask the LLM to provide a summary of the captions in the first round, which we denote as $S$ using the following prompt template: “You are given language descriptions of a video: {$C$}. Please give me a {$N_{w}$} word summary." $N_{w}$ denotes the desired number of words in the summary $S$. Afterward, during the second round of prompting, instead of using the captions $C$, we use the summary $S$ as input for the LLM to select one of the answer candidates. Conceptually, such a prompting scheme is beneficial, as the LLM-generated summary $S$ filters out irrelevant/noisy information from the initial set of captions $C$, making LLM inputs for the subsequent QA process more succinct and cleaner. A detailed illustration of our multi-round prompt is shown in Figure 3.